OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2(x) - 14 sin x + 2 = -5

Atsie (atsie):

https://www.wolframalpha.com/input/?i=7+sin%5E2%28x%29+-+14+sin+x+%2B+2+%3D+-5 I don't know if that will be of any help, but there you go

OpenStudy (freckles):

this is a quadratic in terms of sin(x)

OpenStudy (freckles):

Do you know how to solve the following for u: \[7u^2-14u+2=-5\]

OpenStudy (anonymous):

bring five over and equal it to zero?

OpenStudy (anonymous):

7u^2-14u-3??

OpenStudy (freckles):

but 2+5 is 7 not -3

OpenStudy (freckles):

\[7u^2-14u+2=-5 \\ \text{ add 5 on both sides } \\ 7u^2-14u+2+5=-5+5 \\ 7u^2-14u+7=0\]

OpenStudy (anonymous):

oo ok so now do we combine 7u^2 with 14u?

OpenStudy (freckles):

you can't they aren't like terms

OpenStudy (freckles):

have you ever solved a quadratic before

OpenStudy (freckles):

you can try factoring or use the quadratic formula....or completing the square

OpenStudy (anonymous):

ok lets do quadratic form is A=7u^2 and b=14U and c=7 ?

OpenStudy (freckles):

do you mean a=7, b=-14,c=7 ?

OpenStudy (anonymous):

i got 7=1 for posititive quardatic

OpenStudy (freckles):

I have no clue how you can conclude 7 is 1...

OpenStudy (freckles):

7 is greater than 1 so it cannot be 1

OpenStudy (freckles):

\[au^2+bu+c=0 \\ u=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (freckles):

all you do is enter in the a,b,c into the above formula

OpenStudy (anonymous):

yes i know the rhyme for the quad. o got 1 for positive

OpenStudy (anonymous):

i got 1=1

OpenStudy (freckles):

what did you get for u ?

OpenStudy (anonymous):

i got u equals 1 for both eqautions

OpenStudy (freckles):

\[u=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \\ \text{ we have } a=7,b=-14,c=7 \\ u=\frac{-(-14) \pm \sqrt{(-14)^2-4(7)(7)}}{2(7)} \\ u=\frac{14 \pm \sqrt{196-196}}{14} \\ u=1\]oh so you mean u=1 not 1=1..which doesn't really help us

OpenStudy (freckles):

ok well remember u was in the place of sin(x)

OpenStudy (freckles):

sin(x)=1 when x=?

OpenStudy (anonymous):

1?

OpenStudy (freckles):

sin(1) is not 1

OpenStudy (freckles):

look at the unit circle if you have to for what angles on the unit circle do you have the y-coordinate is 1 on the unit circle

OpenStudy (anonymous):

90 deg

OpenStudy (freckles):

yep so 90 deg+360 deg*k where k is a integer

OpenStudy (freckles):

or you prefer.... pi/2+2pi*k where k is an integer

OpenStudy (anonymous):

is that the answer ?

OpenStudy (freckles):

yep one of those are the answer it depends on the preference of your teacher

OpenStudy (anonymous):

wow thats was alot easier than i expected. when can you not use the quadratic form?

OpenStudy (freckles):

when it can't be written in the form au^2+bu+c=0

OpenStudy (freckles):

if it can be written in the form au^2+bu+c=0 then it is in quadratic form in terms of u whether u be sin(x) or cos(x) or tan(x)...

OpenStudy (anonymous):

ok thank you so much